# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v5

@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) ∗600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) ∗2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python 2024-03 ProCQA Li et al. (2024b) 5.2M C, C++, Java, Python, Ruby, Lisp, JS, C#, Go, Rust, PHP Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) †15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) †13948 2023-06 CMMLU Li et al. (2023b) †11528 2023-09 CodeApex Fu et al. (2023) 250 Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) ‡300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) ‡513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023c) ‡744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) ⋄21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python

### v6

@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) ∗600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) ∗2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python 2024-03 ProCQA Li et al. (2024b) 5.2M C, C++, Java, Python, Ruby, Lisp, JS, C#, Go, Rust, PHP Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) †15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) †13948 2023-06 CMMLU Li et al. (2023b) †11528 2023-09 CodeApex Fu et al. (2023) 250 C++ 2024-01 CRUXEval Gu et al. (2024) 800 Python 2024-05 PythonIO Zhang et al. (2024d) 2650 Python Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) ‡300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) ‡513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023c) ‡744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) ⋄21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python